24. On October 1, 2018, Adidas Corporation purchased copyright to Adidas Jr Company for $234,000. It is estimated that the copyright will have a useful life of 5 years with an estimated salvage value of $24,000. The amount of Amortization Expense recognized for the year 2018would be

2. Using the following information:                          12/31/15

Accounts receivable                                              $525,000

Allowance                                                              (40,000)

Cash realizable value                                            $485,000

During 2016, sales on account were $145,000 and collection on account were $100,000, Also during 2016, the company wrote off $5,000 in uncollectible accounts. An analysis of outstanding receivable accounts at year end indicated that uncollectible accounts should be estimated at $45,000.

Bad debt expense for 2016 is:

3. Sears Corporation bought equipment on December 1, 2018. The equipment cost $360,000 and had an expected salvage value of $60,000. The life of the equipment was estimated to be 5 years. The depreciable cost of the equipment is
Answer #1

1) Amortization expense for 2018 = (234000-24000)/5*3/12 = 10500

2) Bad debt expense = 40000-5000-45000 = 10000

3) Depreciable cost = Cost-Salvage value = 360000-60000 = $300000